# Towards a Theoretical Understanding of Synthetic Data in LLM Post-Training: A Reverse-Bottleneck Perspective

### v2

@@ Abstract
Synthetic data has become a pivotal resource in post-training tasks for large language models (LLMs) due to the scarcity of high-quality, specific data. While various methods have been developed to generate synthetic data, there remains a discernible gap between the practical effects of synthetic data and our theoretical comprehension. To address this challenge, we commence by presenting a detailed modeling of the prevalent synthetic data generation process. Building upon this modeling, we demonstrate that the generalization capability of the post-trained model is critically determined by the information gain derived from the generative model, as analyzed from a novel reverse-bottleneck perspective. Moreover, we introduce the concept of Generalization Gain via Mutual Information (GGMI) and elucidate the relationship between generalization gain and information gain. This analysis serves as a theoretical foundation for synthetic data generation and further highlights its connection with the generalization capability of post-trained models, offering an understanding about the design of synthetic data generation techniques and the optimization of the post-training process. We open source our code at https://github.com/ZyGan1999/Towards-a-Theoretical-Understanding-of-Synthetic-Data-in-LLM-Post-Training.
@@ 2.2 Synthetic Data in LLMs
Large language models (LLMs), a specialized subset of generative models tailored for the text domain, have demonstrated remarkable capabilities in generating high-quality text data. Similar to traditional generative data augmentation, synthetic data produced by these models is increasingly utilized to enhance LLMs, particularly during post-training phases. Given the scarcity of labeled data in specific domains, synthetic data plays a crucial role in boosting the performance of LLMs across a variety of downstream tasks, including text classification (Li et al., 2023), clinical text mining (Tang et al., 2023), and code generation (Tsai et al., 2024). However, unlike classic generative data augmentation, synthetic data within LLMs is typically generated by the language models themselves and often predominates the training data in post-training stages. This predominance stems from the high-quality demands of synthetic data in LLM contexts, which necessitates alignment with human intent. Efforts to enhance the quality of synthetic data in LLMs have included integrating methodologies such as active learning (Wagner et al., 2024) and reinforcement learning (Setlur et al., 2024). Despite these advancements, the theoretical understanding of how synthetic data influences the learning process in LLMs remains limited. Key questions persist regarding the mechanisms through which synthetic data impacts LLM training and the optimal strategies for designing synthetic data to maximize LLM performance (Long et al., 2024). Addressing these questions is essential for furthering our comprehension and utilization of synthetic data in enhancing large language model efficacy.
@@ 3.1 Notations & Experimental Setup
Let $S_{\text{anchor}}$ represent the real data utilized for generation, and $S_{\text{gen}}$ denote the synthetically generated data. The LLM employed in the generation process is designated as $M$, with the input prompt labeled as $p$. The distribution of the post-training target task $\mathcal{T}$ is referred to as $\mathcal{D}$, while the output distribution of the LLM is denoted by $\mathcal{D}_{M}$. Additionally, the distribution corresponding to the synthetic data is represented as $\mathcal{D}_{\text{gen}}$. The generalization error associated with the under-aligned LLM $\pi$ on the synthetic data $S_{\text{gen}}$ is expressed as $\operatorname{Err}(\pi^{S_{\text{gen}}})$, and the generalization error related to the anchor data is indicated by $\operatorname{Err}(\pi^{S_{\text{anchor}}})$. We define $H(\cdot)$ as the entropy of a random variable, $I(\cdot,\cdot)$ as the mutual information between two random variables, $D_{\text{KL}}$ as the Kullback-Leibler divergence, and $D_{\text{TV}}$ as the total variation distance. The detailed definitions are listed in Appendix A.
@@ 3.2 Modeling Synthetic Data Generation
The prompt $p$ is intricately linked to the downstream task $\mathcal{T}$ and the anchor data $S_{\text{anchor}}$. We postulate that $S_{\text{anchor}}$ forms the core of the prompt $p$, upon which a task-specific transformation function $\phi_{\mathcal{T}}$ is applied. Consequently, the prompt $p$ can be mathematically modeled as $p=\phi_{\mathcal{T}}(S_{\text{anchor}})$, where $\phi_{\mathcal{T}}$ is a reversible function that maps the anchor data to the prompt, consists of all the task-relevant transformation, like the template and other customized settings for more faithful and diverse generation.
@@
To provide a clearer visualization, we simulate the distribution relationships using GMMs, the result is depicted in Figure 2. The anchor data is represented by blue dots, and their distribution is illustrated by blue ellipses. In contrast, synthetic data is represented by orange. The distributions of $S_{\text{gen}}$ are visualized as an effort to encompass the distributions of $S_{\text{anchor}}$. However, since $S_{\text{gen}}$ is derived from the model $M$, which incorporates more complex distribution components, the distribution of $S_{\text{gen}}$ not only attempts to mirror $S_{\text{anchor}}$ but also extends beyond, covering broader areas.
@@ 4 Main Result
In preceding sections, we established a comprehensive modeling for synthetic data generation and elucidated the connection between this process and the generalization error as delineated in Lemma 3.1. This section delves deeper into the implications of the synthetic data generation process on the generalization capabilities.
@@ 4.1 Information Gain & Reverse-Bottleneck
Following this framework, we proceed to introduce the concept of information gain within the context of the synthetic data generation process.
@@ 4.3 Generalization Gain with Synthetic Data
A larger upper bound for the GGMI signifies greater potential generalization benefits when utilizing synthetic data. To elucidate the impact of synthetic data on model generalization, we isolate the influence of $W^{{}^{\prime}}$ and establish that the GGMI can be effectively bounded. This relationship is formalized in Theorem 4.10, which demonstrates that GGMI is constrained by $\Delta I$—the measure of information gain—and various entropy components solely associated with $W$.
@@
As emphasized in (Long et al., 2024), the generation of synthetic data typically focuses on two primary objectives: faithfulness and diversity. These objectives are associated with $H(S_{\text{anchor}}|W)$ and $\Delta I$, respectively. Specifically, $H(S_{\text{anchor}}|W)$, which quantifies the relevance between $W$ and $S_{\text{anchor}}$, as presented in Theorem 4.10, encourages the model to align with the information derived from the original real data. This alignment is crucial for achieving the objective of faithfulness. In addition, $\Delta I$ serves as a measurement of the additional information introduced by the generative model $M$. Given that $M$ is typically pre-trained on a more extensive dataset, $\Delta I$ in Theorem 4.10 promotes the objective of diversity by facilitating greater information gain from $M$.
@@ 4.4 Verification with GMM Simulation
To control the variables outlined in Theorem 4.10, we adjust the number of components in the GMM $M$ and the ground-truth GMM $G$. The result is illustrated in Figure 4. Generally, increasing $J$ facilitates the scaling of $\Delta I$, resulting in a larger upper bound for GGMI. In contrast, larger $K$ amplifies the influence of anchor data within the post-training target distribution, thereby increasing the $H(S_{\text{anchor}}|W)$ term and tightening the upper bound of GGMI. Additionally, while an increase in $L$ enhances $H(S_{\text{gen}}|W)$ due to the introduction of greater diversity, it concurrently leads to a reduction in $\Delta H$. As a result, we observe a trade-off manifested as a decrease in the KL Gap in our simulation outcomes.
@@ 5 Conclusion
In this paper, we have conducted a detailed analysis of synthetic data utilization in post-training large language models (LLMs). We present a comprehensive modeling of the current synthetic data generation process, focusing on its distributional aspects, which further connects the generalization capabilities of post-trained models. We introduce a novel reverse-bottleneck framework, allowing us to derive a measurable upper bound on generalization errors. Our analysis reveals that the pivotal constraint on generalization ability is influenced by the information gain from the generative model $M$. Additionally, we present the Generalization Gain via Mutual Information (GGMI), showing that larger information gains enhance the generalization capability of post-trained models. We emphasize the importance of balancing faithfulness and diversity during post-training stages, providing a theoretical foundation for existing methodologies.
Unfortunately, due to limitations in computational resources, we are unable to validate our findings within real-world LLM settings. Looking ahead, future research should focus on developing adaptive models that respond to the evolving characteristics of synthetic data. This includes enhancing generative models and fine-tuning parameters for specific learning scenarios, as well as exploring various generative models to better replicate real-world data complexities while improving model performance.
Appendix A Definition of Notations

### v3

@@ Abstract
Synthetic data has become a pivotal resource in post-training tasks for large language models (LLMs) due to the scarcity of high-quality, specific data. While various methods have been developed to generate synthetic data, there remains a discernible gap between the practical effects of synthetic data and our theoretical comprehension. To address this challenge, we commence by presenting a detailed modeling of the prevalent synthetic data generation process. Building upon this modeling, we demonstrate that the generalization capability of the post-trained model is critically determined by the information gain derived from the generative model, as analyzed from a novel reverse-bottleneck perspective. Moreover, we introduce the concept of Generalization Gain via Mutual Information (GGMI) and elucidate the relationship between generalization gain and information gain. This analysis serves as a theoretical foundation for synthetic data generation and further highlights its connection with the generalization capability of post-trained models, offering an understanding about the design of synthetic data generation techniques and the optimization of the post-training process. We open-source our code at https://github.com/ZyGan1999/Towards-a-Theoretical-Understanding-of-Synthetic-Data-in-LLM-Post-Training.
@@ 2.2 Synthetic Data in LLMs
Similar to traditional generative data augmentation, synthetic data produced by LLMs is increasingly utilized to enhance post-training phases. Given the scarcity of labeled data in specific domains, synthetic data plays a crucial role in boosting the performance of LLMs across a variety of downstream tasks, including text classification (Li et al., 2023), clinical text mining (Tang et al., 2023), and code generation (Tsai et al., 2024). However, unlike classic generative data augmentation, synthetic data within LLMs is typically generated by the language models themselves and often predominates the training data in post-training stages. This predominance stems from the high-quality demands of synthetic data in LLM contexts, which necessitates alignment with human intent. Efforts to enhance the quality of synthetic data in LLMs have included integrating methodologies such as active learning (Wagner et al., 2024) and reinforcement learning (Setlur et al., 2024). Despite these advancements, the theoretical understanding of how synthetic data influences the learning process in LLMs remains limited. Key questions persist regarding the mechanisms through which synthetic data impacts LLM training and the optimal strategies for designing synthetic data to maximize LLM performance (Long et al., 2024). Addressing these questions is essential for furthering our comprehension and utilization of synthetic data in enhancing large language model efficacy.
@@ 3.1 Notations & Experimental Setup
Let $S_{\text{anchor}}$ represent the real data utilized for generation, and $S_{\text{gen}}$ denote the synthetically generated data. The LLM employed in the generation process is designated as $M$, with the input prompt labeled as $p$. The distribution of the post-training target task $\mathcal{T}$ is referred to as $\mathcal{D}$, while the output distribution of the LLM is denoted by $\mathcal{D}_{M}$. Additionally, the distribution corresponding to the synthetic data is represented as $\mathcal{D}_{\text{gen}}$. The generalization error associated with the under-aligned LLM $\pi$ on the synthetic data $S_{\text{gen}}$ is expressed as $\operatorname{Err}(\pi^{S_{\text{gen}}})$, and the generalization error related to the anchor data is indicated by $\operatorname{Err}(\pi^{S_{\text{anchor}}})$. We define $H(\cdot)$ as the entropy of a random variable, $I(\cdot,\cdot)$ as the mutual information between two random variables, $D_{\text{KL}}$ as the Kullback-Leibler divergence, and $D_{\text{TV}}$ as the total variation distance. The detailed definitions are listed in Appendix A.1.
@@ 3.2 Modeling Synthetic Data Generation
The prompt $p$ is intricately linked to the downstream task $\mathcal{T}$ and the anchor data $S_{\text{anchor}}$. We postulate that $S_{\text{anchor}}$ forms the core of the prompt $p$, upon which a task-specific transformation function $\phi_{\mathcal{T}}$ is applied. Consequently, the prompt $p$ can be mathematically modeled as $p=\phi_{\mathcal{T}}(S_{\text{anchor}})$, where $\phi_{\mathcal{T}}$ is a function that maps the anchor data to the prompt, consists of all the task-relevant transformation, like the template and other customized settings for more faithful and diverse generation.
@@
To provide a clearer visualization, we simulate the distribution relationships using GMMs, the result is depicted in Figure 2. The distributions of $S_{\text{gen}}$ are visualized as an effort to encompass the distributions of $S_{\text{anchor}}$. However, since $S_{\text{gen}}$ is derived from the model $M$, which incorporates more complex distribution components, the distribution of $S_{\text{gen}}$ not only attempts to mirror $S_{\text{anchor}}$ but also extends beyond, covering broader areas.
@@ 4 Main Result
In this section, we delves deeper into the implications of the synthetic data generation process on the generalization capabilities.
@@ 4.1 Information Gain & Reverse-Bottleneck
With the synthetic factors established, we posit that the synthetic data $S_{\text{gen}}$ is primarily governed by two distinct factors: $e_{p}$ and $e_{M}$, which are actually assumed random variables related to the prompt $p$ and the LLM $M$ respectively. Following this framework, we proceed to introduce the concept of information gain within the context of the synthetic data generation process.
@@ 4.3 Generalization Gain with Synthetic Data
A larger upper bound for the GGMI signifies greater potential generalization benefits when utilizing synthetic data. To elucidate the impact of synthetic data on model generalization, we isolate the influence of $W^{{}^{\prime}}$ and establish that the GGMI can be effectively bounded.
@@
As emphasized in (Long et al., 2024), the generation of synthetic data typically focuses on two primary objectives: faithfulness and diversity. These objectives are associated with $\Delta H$ and $\Delta I$, respectively. Specifically, $\Delta H$, which quantifies the entropy decrease during synthetic data generation, as presented in Theorem 4.10, encourages the model to eliminate uncertainty during synthetic data generation, thereby enhancing the faithfulness of the synthetic data. In addition, $\Delta I$ serves as a measurement of the additional information introduced by the generative model $M$. Given that $M$ is typically pre-trained on a more extensive dataset, $\Delta I$ in Theorem 4.10 promotes the objective of diversity by facilitating greater information gain from $M$.
@@ 4.4 Verification with GMM Simulation
To control the variables outlined in Theorem 4.10, we adjust the number of components in the GMM $M$ and the ground-truth GMM $G$. The result is illustrated in Figure 4. Generally, increasing $J$ facilitates the scaling of $\Delta I$, resulting in a larger upper bound for GGMI. In contrast, larger $K$ amplifies the influence of anchor data within the post-training target distribution, thereby increasing the $H(S_{\text{anchor}}|W)$ term and tightening the upper bound of GGMI. Additionally, while an increase in $L$ enhances $H(S_{\text{gen}}|W)$, it concurrently leads to a reduction in $\Delta H$. As a result, we observe a trade-off manifested as a decrease in the KL Gap in our simulation outcomes.
@@ 5 Conclusion
In this paper, we have conducted a detailed analysis of synthetic data utilization in post-training large language models (LLMs). We present a comprehensive modeling of the current synthetic data generation process, focusing on its distributional aspects, which further connects the generalization capabilities of post-trained models. We introduce a novel reverse-bottleneck framework, allowing us to derive a measurable upper bound on generalization errors. Our analysis reveals that the pivotal constraint on generalization ability is influenced by the information gain from the generative model $M$. Additionally, we present the Generalization Gain via Mutual Information (GGMI), showing that larger information gains enhance the generalization capability of post-trained models. We emphasize the importance of balancing faithfulness and diversity during post-training stages, providing a theoretical foundation for existing methodologies. Unfortunately, due to limitations in computational resources, we are unable to validate our findings within real-world LLM settings. Looking ahead, future research should focus on developing adaptive models that respond to the evolving characteristics of synthetic data. This includes enhancing generative models and fine-tuning parameters for specific learning scenarios, as well as exploring various generative models to better replicate real-world data complexities while improving model performance.
Acknowledgments
This research was supported by National Natural Science Foundation of China (No.62476277), National Key Research and Development Program of China (No.2024YFE0203200), CCF-ALIMAMA TECH Kangaroo Fund (No.CCF-ALIMAMA OF 2024008), and Huawei-Renmin University joint program on Information Retrieval. We also acknowledge the support provided by the fund for building worldclass universities (disciplines) of Renmin University of China and by the funds from Beijing Key Laboratory of Big Data Management and Analysis Methods, Gaoling School of Artificial Intelligence, Renmin University of China, from Engineering Research Center of Next-Generation Intelligent Search and Recommendation, Ministry of Education, from Intelligent Social Governance Interdisciplinary Platform, Major Innovation & Planning Interdisciplinary Platform for the “DoubleFirst Class” Initiative, Renmin University of China, from Public Policy and Decision-making Research Lab of Renmin University of China, and from Public Computing Cloud, Renmin University of China.
Appendix A Definition and Introduction about Infomation Bottleneck Theory
A.1 Definition of Notations
@@
A.2 The Information Bottleneck Theory
The information bottleneck (IB) theory is a theoretical construct designed to elucidate the learning processes within neural networks. In essence, for a given Markov chain $X\rightarrow Z\rightarrow Y$, the IB theory aims to optimize the learning process by maximizing the mutual information between $Y$ and $Z$ while minimizing the mutual information between $X$ and $Z$. The optimization objective in IB theory is generally expressed as:
$\mathcal{L}\left[p\left(z|x\right)\right]=I(Z,X)-\beta I(Z,Y).$ (11)
Originally developed within the context of information compression, IB theory has been widely adopted across various deep learning fields, and further research has explored generalization error upper bounds that incorporate mutual information (Russo & Zou, 2019; Xu & Raginsky, 2017). These studies have established a connection between the generalization capabilities of deep neural networks (DNNs) and IB theory (Alquier et al., 2024). A representative formulation of a generalization error upper bound from the perspective of mutual information is as follows:
$\operatorname{genErr}\leq\sqrt{\frac{2\sigma^{2}}{n}I(S,W)},$ (12)
where $S$ and $W$ are training data and model parameters respectively, with the assumption that the loss is $\sigma$-subgaussian. This type of bound suggests that the generalization error is intrinsically limited by the relevance between the training data and the learned model parameters.
@@ Appendix C Details of Synthetic Data Generation Modeling
Appendix G Experiments: Exploring Better Synthetic Data in Practice
To further investigate the process of synthetic data generation in real-world settings, we conduct experiments to evaluate the quality of synthetic data produced under different conditions and aim to identify the factors that contribute to its effectiveness in enhancing model performance.
The experiments follow the same setup as described in the main text, with the synthetic data $S_{\text{gen}}$ generated from a generative model $M$ prompted by $p$. We utilize a standard in-context learning (ICL) framework to determine $p$ using anchor data $S_{\text{anchor}}$, and we then evaluate the performance of the model trained on the synthetic data. Additionally, we estimate key components from our theoretical analysis in the main text, including information gain $\Delta I$ and the entropy of the synthetic data $H(S_{\text{gen}})$.
In the remainder of this section, we commence by introducing the experimental setup and the evaluation metrics. We then present the results of the synthetic data generation, focusing on the performance of the model trained on synthetic data to assess its quality. Furthermore, we estimate the key components from our theoretical analysis and analyze the factors that contribute to the effectiveness of synthetic data in improving model performance. Finally, we provide a brief conclusion and discuss potential principles for generating higher-quality synthetic data in practice.
G.1 Experimental Setup
We conducted experiments to evaluate the effectiveness of synthetic data generated by a generative model $M$ prompted by $p$ in enhancing model performance. Our experimental setup follows the synthetic data utilization process described in the main text, including selecting benchmark dataset, determining prompt $p$, generating synthetic data $S_{\text{gen}}$, training the model on the synthetic data, and evaluating the trained model.
G.1.1 Benchmark Dataset
The benchmark dataset is utilized to sample $S_{\text{anchor}}$. Specifically, we adopt Dolly-15K (Conover et al., 2023) as our benchmark dataset, which contains 15,000 lines of text data designed for instruction-following tasks. We split the benchmark dataset into training and testing sets with a ratio of 8:2, and $S_{\text{anchor}}$ is sampled from the training set. For each data instance, we retain the keys “instruction”, “context” and “response” and combine them using the following template.
G.1.2 Determining Prompt $p$
Consistent with the methodology described in the main text, we employ a standard In-Context Learning (ICL) framework to determine the prompt $p$. Specifically, $p=E(S_{\text{anchor}})$, where $E$ is a predefined template for the prompt. We follow the settings of Alpaca (Taori et al., 2023) and modify the template to better suit the benchmark dataset used in our experiments. The modified template is as follows:
You are asked to come up with a set of 20 diverse task instructions. These task instructions will be given to a language model and we will evaluate the language model for completing the instructions. Here are the requirements: 1. Try not to repeat the verb for each instruction to maximize diversity. 2. The language used for the instruction also should be diverse. For example, you should combine questions with imperative instrucitons. 3. The type of instructions should be diverse. The list should include diverse types of tasks like open-ended generation, classification, editing, etc. 4. The language model should be able to complete the instruction. For example, do not ask the assistant to create any visual or audio output. For another example, do not ask the assistant to wake you up at 5pm or set a reminder because it cannot perform any action. 5. The instructions should be in English. 6. The instructions should be 1 to 2 sentences long. Either an imperative sentence or a question is permitted. 7. You should generate an appropriate input to the instruction. The input field should contain a specific example provided for the instruction. It should involve realistic data and should not contain simple placeholders. The input should provide substantial content to make the instruction challenging but should ideally not exceed 100 words. 8. Not all instructions require input. For example, when a instruction asks about some general information, ”what is the highest peak in the world”, it is not necssary to provide a specific context. In this case, we simply put “noinput” in the input field. 9. The output should be an appropriate response to the instruction and the input. Make sure the output is less than 100 words. Your output should consist of 3 parts: instruction, context and reference response. “Instruction” is the task instruction which the language model should complete. “Context” is information related to the instruction, if don’t need, you can set it as empty. “Reference response” is the correct answer to the instruction your recommend. Your output must be in the following json form like: {“instruction”: [the instruction you generate], “context”: [the context you generate], “reference response”: [the reference response you generate]} Here are some examples you should emulate: {anchor data} List of 20 tasks:”’
We then sample $S_{\text{anchor}}$ from the benchmark dataset and populate the “{anchor data}” placeholders in the prompt template with these samples. This completes the process of determining the prompt $p$.
G.1.3 Generating Synthetic Data $S_{\text{gen}}$
After determining the prompt $p$, we generate synthetic data $S_{\text{gen}}$ by prompting the generative model $M$ with $p$. In our experiments, we primarily utilize GPT-4o (OpenAI, 2024) as the generative model $M$. Additionally, wo also employ the latest Llama 3.2 models (Meta, 2024) including Llama-3.2-1B-Instruct and Llama-3.2-3B-Instruct for comparison experiments.
G.1.4 Training on Synthetic Data
We fine-tune a GPT-2 (Radford et al., 2019) model using both the synthetic data $S_{\text{gen}}$ generated by the generative model $M$ and the training set $T$ of the benchmark dataset. The training procedure follows the standard instruction tuning process, where we fine-tune the model on the synthetic data for a fixed 20 epochs.
G.1.5 Evaluating Fine-tuned Model
We assess the performance of the fine-tuned model on the testing set of the benchmark dataset. Following the evaluation procedure of  Zheng et al. (2023c), we evaluate the model’s ability by rating the generated responses using a LLM. To better align our evaluation with our datasets, we modify the original evaluation prompt to ensure that the judge LLM compares the output with the ground-truth answer. The evaluation prompt we adopt is as follows:
Please act as an impartial judge and evaluate the quality of the response provided by an AI assistant to the user question displayed below. You are provided with 4 parts of the text, [Question] is the question asked by the user, [Context] is information related to the question, [Reference Answer] is the correct answer for your reference, and Assistant’s Answer which is surrounded by [The Start of Assistant’s Answer] and [The End of Assistant’s Answer] is the answer given by the assistant. Your evaluation should consider factors such as the helpfulness, relevance, accuracy, depth, creativity, and level of detail of the response. Begin your evaluation by providing a short explanation. Be as objective as possible. After providing your explanation, you must rate the Assistant’s Answer on a scale of 1 to 10 by strictly following this format: “[[rating]]”, for example: “Rating: [[5]]”. [Question] {instruction} [Context] {context} [Reference Answer] {reference response} [The Start of Assistant’s Answer] {generated response} [The End of Assistant’s Answer]
We then populate the placeholders “{instruction}”, “{context}”, “{reference response}”, and “{generated response}” in the evaluation prompt with the corresponding text. We adopt Llama-3.1-Nemotron-70B-Instruct-HF (Wang et al., 2024) as the judge LLM and extract the ratings from its output. The final rating is averaged over the testing set to evaluate the performance of the fine-tuned model.
G.2 Synthetic Data Quality
We assess the quality of synthetic data generated by the generative model $M$ prompted by $p$ in terms of its effectiveness in enhancing model performance. Specifically, we utilize GPT-4o as $M$ to generate synthetic data with varying numbers of instances in ICL, corresponding to different sizes of $S_{\text{anchor}}$, denoted as 3-ins, 10-ins, and 20-ins, respectively. We then fine-tune a GPT-2 model on both the synthetic data and the training set of the benchmark dataset. The performance of the fine-tuned model on the testing set is used as a measure of the quality of the synthetic data. For better presentation, we apply a softmax function to normalize the ratings. The results are shown in Table 1.
Synthetic Data Fine-Tuned
Base 3-ins 10-ins 20-ins Real Data Fine-Tuned
Rating 0.1409 0.1863 0.1965 0.2015 0.2745
Table 1: Average ratings of the fine-tuned model on the testing set. The ratings were normalized using a softmax function. The synthetic data were generated by GPT-4o with varying numbers of instances in In-Context Learning (ICL) (denoted as $x$-ins). The unfine-tuned base model (Base) and the model fine-tuned on real data are marked gray.
The results demonstrate that the synthetic data effectively enhances the performance of the fine-tuned model, with the rating positively correlated with the number of instances in ICL. This finding indicates that appropriately increasing the number of instances in ICL can improve the quality of the synthetic data. This phenomenon may be attributed to the fact that increasing the number of instances in the ICL prompts provides the generative model with a richer and more diverse context. This enhanced context allows the model to capture a broader range of patterns present in the anchor data, thereby generating synthetic data with richer content.
G.3 Estimating Theoretical Components
Building upon the results of synthetic data quality, we further estimate the key components from our theoretical analysis, including information gain $\Delta I$ and the entropy of the synthetic data $H(S_{\text{gen}})$. We aim to analyze the factors that contribute to improving the quality of synthetic data.
G.3.1 Estimating Information Gain
Given the definition of information gain $\Delta I$ in Definition 4.2, it is difficult to directly estimate $\Delta I$ in practice. However, it is possible to estimate $I(T,S_{\text{gen}})$, the mutual information between the synthetic data $S_{\text{gen}}$ and the training set $T$ of the benchmark dataset where $S_{\text{anchor}}$ is sampled. Since the crucial part of prompt $p$ is $S_{\text{anchor}}$, $I(T,S_{\text{gen}})$ has a negative correlation with $\Delta I$ to a certain extent.
To measure $I(T,S_{\text{gen}})$, we follow the setting of existing works (Qian et al., 2024; Ma et al., 2020) and utilize HSIC (Gretton et al., 2005) as an estimator. The result is shown in Table 2.
$S_{\text{gen}}$ 3-ins 10-ins 20-ins
HSIC w/ $T$ (×$\times 10^{-3}$) 7.8703 7.8668 7.8502
Table 2: The HSIC value between the synthetic data $S_{\text{gen}}$ and training set $T$ for different numbers of instances in ICL setting.
It is supurising that more instances doesn’t increase the HSIC value, but even lead to a lower HSIC value, indicating reduced mutual information between the synthetic data and the training set. This phenomenon suggests that enlarging the sizes of $S_{\text{anchor}}$ does not significantly increase the dependency between the synthetic data and the training set, and may even enhance the diversity of the synthetic data. This may be attributed to the fact that when a LLM with a wide range of knowledge is employed as $M$, it leverages its broad understanding to generate synthetic data that is less reliant on the specific instances in $S_{\text{anchor}}$. As the number of instances in the ICL setup increases, the LLM interprets this as a richer and more varied context, thereby increasing the output diversity instead.
A smaller HSIC value indicates a lower mutual information between the synthetic data and the training set, which leads to a larger information gain $\Delta I$. With Theorem 4.7 and Theorem 4.10, this guarantees a tighter upper bound of the generalization error and higher GGMI, which contributes to the quality of synthetic data and increase the generalization capabilities.
G.3.2 Estimating Entropy of Synthetic Data
As another important component in Theorem 4.10, $H(S_{\text{gen}})$ is crutial for determining the value of $\Delta H$. We use semantic entropy (Farquhar et al., 2024) as an estimator to measure the entropy of the dataset and estimate the value of $H(S_{\text{gen}})$. The result is shown in Table 3.
$S_{\text{gen}}$ 3-ins 10-ins 20-ins
Semantic Entropy 1.0739 1.0503 1.0005
Table 3: The semantic entropy of the synthetic data $S_{\text{gen}}$ in different numbers for instances in ICL setting.
The results indicate that the semantic entropy of the synthetic data $S_{\text{gen}}$ is also negatively correlated with the number of instances in ICL. This suggests that increasing the sizes of $S_{\text{anchor}}$ when utilizing LLM as generative model $M$ can help reduce the entropy of the synthetic data. This reduction in entropy may be attributed to the richer and more varied context provided by a larger $S_{\text{anchor}}$, which enables $M$ to generate more accurate and informative synthetic data, thereby increasing the faithfulness of the synthetic data.
A smaller semantic entropy indicates a lower entropy of the synthetic data $S_{\text{gen}}$, which leads to a larger $\Delta H$. With Theorem 4.10, this benifts increasing the upper bound of GGMI, and contributes to the generalization capabilities of the model trained on the synthetic data.
G.4 Estimating on Different Model Architectures
To further investigate the impact of different model architectures and parameters on the quality of synthetic data, we conduct experiments to evaluate the HSIC value and semantic entropy of the synthetic data $S_{\text{gen}}$ generated by different models. Due to computational resource limitations, we utilized GPT-4o, Llama-3.2-3B-Instruct, and Llama-3.2-1B-Instruct as the generative model $M$ to generate synthetic data with 3 instances in ICL setting. The results are presented in Table 4.
GPT-4o Llama-3.2-3B-Instruct Llama-3.2-1B-Instruct
HSIC w/ $T$ (×$\times 10^{-3}$) 7.8703 11.4306 /
Semantic Entropy 1.0739 2.9427 /
Table 4: The HSIC value and semantic entropy of the synthetic data $S_{\text{gen}}$ generated using different model architectures. All the synthetic data are generated with 3 instances in ICL setting. Note that the Llama-3.2-1B-Instruct model did not adhere to the format requirements and thus failed to produce meaningful synthetic data.
Note that under the prompt determined in the experimental setups, the Llama-3.2-1B-Instruct model did not adhere to the format requirements and failed to produce meaningful synthetic data. Consequently, the estimators are not available for this model. This observation underscores a fundamental premise that the generative model $M$ must possess sufficient instruction-following capabilities to generate synthetic data that can be effectively utilized to enhance model performance.
On the other hand, although Llama-3.2-3B-Instruct produced usable synthetic data, its quality was insufficient for fine-tuning GPT-2, and the HSIC value and semantic entropy were significantly higher than those of GPT-4o. This may be attributed to the smaller model size of Llama-3.2-3B-Instruct compared to GPT-4o, resulting in a diminished capacity to generate synthetic data that is both faithful to and diverse from the anchor data. For instance, we provide some examples of the synthetic data generated by Llama-3.2-3B-Instruct in the following as a case study:
Instructions generated by Llama-3.2-3B-Instruct: “instruction”: “Explain the concept of blockchain in simple terms.” “instruction”: “Explain the concept of artificial intelligence in simple terms.” “instruction”: “Explain the concept of climate change in simple terms.” $\cdots$ “instruction”: “Identify the type of music genre: classical or jazz: ’Moonlight Sonata’ or ’Take Five’” “instruction”: “Identify the type of literary device used in the following sentence: ’The eyes that fixed upon her were like two bright stars in the night sky.’” “instruction”: “Identify the type of music instrument: string or woodwind: ’Violin’ or ’Flute’” $\cdots$ “instruction”: “Write a short story about a character who discovers a hidden world within their reflection.” “instruction”: “Write a review of the movie ’The Shawshank Redemption’.” “instruction”: “Write a poem about the beauty of nature.”
The examples demonstrate that the synthetic data generated by Llama-3.2-3B-Instruct is highly homogeneous, even within a single generation cycle. Moreover, it is highly dependent on the specific instances in the anchor data, leading to a higher HSIC value. Furthermore, although the synthetic data lacks diversity in form, the semantic entropy remains high. This indicates that the generated synthetic data lacks sufficient faithfulness. Collectively, these factors contribute to the poor quality of the synthetic data produced by Llama-3.2-3B-Instruct.
G.5 Conclusion
Building upon the experiments, we can derive some brief conclusions about how to guarantee the synthetic data quality and estimate the key factors in real-world LLM practice.
The quality of synthetic data is mainly reflected in two aspects: diversity and faithfulness. Diversity makes the synthetic data contain richer contents and thus increase the information gain $\Delta I$. With our theoretical analysis, this will benifit the generalization ability of the model post-trained on synthetic data. Faithfulness makes the synthetic data semantically continuous, and thus decrease the entropy of the synthetic data $S_{\text{gen}}$, which also strenghten the generalization capabilities.
In practice, the diversity and the faithfulness can be estimated by HSIC value and the semantic entropy, respectively, as demonstrated in the experimental settings of this section. It is also important to highlight that employing a generative model with stronger instruction-following capabilities and more diverse knowledge can enhance the quality of synthetic data in both aspects.